At the beginning of 2019, Ace Company had the following portfolio of investments in available-for-sale debt securities (all of which were acquired at par value):

Security

Cost

1/1/19 Fair Value

A $15,000 $19,000
B 23,000 22,000
Totals $38,000 $41,000

During 2019, the following transactions occurred:

Transactions:
May 3 Purchased C debt securities at their par value for $45,000.
July 1 Sold all of the A securities for $19,000 plus interest of $1,000.
Dec. 31 Received interest of $1,000 on the B and C securities. Additionally the following information was available:

Security

12/31/19 Fair Value

B $26,000
C 47,000

Required:

1. Prepare journal entries to record the preceding information.
2. What is the balance in the Unrealized Holding Gain/Loss account on December 31, 2019?
3. Next Level What justification does the FASB give for its treatment of unrealized holding gains and losses for available-for-sale securities?
Answer #1

1) Journal Entries

Date

General Journal

Debit

Credit

May.3, 2019

Available for sale debt securities

$45,000

Cash

$45,000

(Purchased debt securities)

July.1, 2019

Cash

$20,000

Unrealized Gain - OCI

$4,000

Available for sale debt securities (A)

$19,000

Interest Income

$1,000

Realized Gain on Sale of Debt Security A

$4000

Dec.31, 2019

Cash

$1000

Interest Income

$1000

(Interest received on B and C debt securities)

Dec.31, 2019

Available for sale debt securities (B)

$5,000

Available for sale debt securities (C )

$3,000

   Unrealized Gain (loss) - OCI

$8,000

(Fair value adjustment as on Dec31, 2018)

(Unrealized Gain (B) = 34000 - 29,000 = $5,000)

(Unrealized Gain (C ) = 53,000 - 50,000 = $3,000)

Note 1 --- Unrealized Gain / (Loss) as on 1/1/19

Cost

Fair Value 1/1/19

Unrealized Gain (Loss)

Security A

$15,000

$19,000

$4000

Security B

$23000

$22000

-$1,000

2) Unrealized Gain (loss) balance as on Dec.31, 2019

Beginning balance of Unrealized Gain / (Loss) of Security B

($1,000)

Plus: Fair Value adjustment as on Dec.31, 2019

-- Unrealized Gain on available for sale securities as on Dec31, 2019

    Security B (26000 - 22,000)

$4000

    Security C (47,000 – 45,000)

$2,000

Unrealized Gain Account balance as on Dec 31, 2019

$6,000

3) Debt securities that are bought and held principally for the purpose of selling them in the near term are classified as available for sale securities and reported at fair value, with unrealized gain and losses included in earnings.

The unrealized gain will not be actually realized until the company actually sells the stock and collects the cash.